This is from my nephews Home School course. Thank you for your help in advance. "The capacities of two trucks are 3 tons and 4 tons respectively. If the smaller truck makes 18 more trips than the larger, it can deliver 12 more tons of freight than the larger. How many trips does each truck make?"
Answer by ptaylor(1894)


Let x=number of trips of the larger truck
Then x+18=number of trips of the smaller truck
3(x+18)=amount of freight delivered by the smaller truck
4x=amount of freight delivered by the larger truck
Now we are told that:
3(x+18)=4x+12 get rid of parens (distributive law)
3x+54=4x+12 subtract 4x and also 54 from each side
3x+54-54-4x=4x-4x+12-54 collect like terms
-x=-42 divide both sides by -1
x=42 trips-----------------------number of trips of the larger truck
x+18=42+18=60-------------------------number of trips of the smaller truck
CK
3*60=4*42+12
180=168+12
180=180

Hope this helps---ptaylor